4. (06 pts) The HP ProBook 430 G5 is sold with various configurations, including one with an Intel? CoreTM i3-7020U 2.3 GHz dual-core CPU. Since it has two processors, the CPU can execute two instructions per cycle or 4.6 billion instructions per second a) Assuming maximum parallelism, how long would it take for the G5 to run a program requiring the execution of 100 million instructions? b) Assuming maximum parallelism, how long would it take for the G5 to run a program requiring the execution of 500 billion instructions? c) Redo parts a) and b) assuming the CPU could only execute instructions simultaneously 30% of the time.
Answer #1

4a) 100 * 10^6 / 4.6 * 10^9 = 0.0217 sec

b) 500 * 10^9 / 4.6 * 10^9 = 108.69 sec

c) If parallelism is 30%, then the effective number of instructions per second = 0.3*4.6 + 0.7*2.3 = 2.99 billion instructions per second

100 million

100*10^6/2.99*10^9= 0.0334 sec

500 billion

500*10^9/2.99*10^9 = 167.22 sec